# Braneworld Black Bounce to Transversable Wormhole

### v1

@@ Abstract
We provide a $5D$ generalization of the Hawking et al. black string in a Randall Sundrum setup. The embedded $4D$ geometry corresponds to the Simpson-Visser (SV) spacetime, which can represent a traversable wormhole, a one-way wormhole, or a regular black hole. To achieve this, we linearly deform both the bulk geometry and the bulk matter distribution with respect to a coupling constant. These deformations induce a transition from a $5D$ vacuum AdS distribution to an anisotropic distribution. The latter results in the induced geometry on the brane transitioning from a singular Schwarzschild spacetime to a regularized SV spacetime. Since there are no sources or matter fields on the brane, we can assert that the induced SV geometry on the brane arises from the influence of geometrical and matter deformations in the bulk. Thus, the central singularity is suppressed. We determine the cases where the energy conditions are either satisfied or violated. Our spacetime is asymptotically radial AdS, which is intriguing given the absence of a global AdS box that would prevent instability under larger wavelength perturbations. Therefore, it is no longer appropriate to claim that instability exists for very small perturbations near the AdS horizon.
@@ I Introduction
The RS model involves the embedding of $4D$ flat branes, i.e., whose geometry corresponds to Minkowski spacetime. Since its inception, various generalizations of the Randall-Sundrum model, featuring different types of four-dimensional geometries including black holes and wormholes, have emerged. The approach to studying these generalizations is primarily based on two main approximations: one originates from the brane, and the other from the bulk. In the former, the four-dimensional black holes or wormholes are constructed by studying the equations of motion induced on the brane Shiromizu et al. (2000). Some examples in references Molina and Neves (2012); Neves (2015); Neves and Molina (2012); Casadio et al. (2015). However, this methodology does not fully allow for testing the influence of bulk geometry on the physical and geometric behavior of the four-dimensional geometry. Regarding the latter, the second approach constructs spacetimes directly from the bulk, and the brane equations are derived from there. See Kanti et al. (2002); Creek et al. (2006). The approach embraced in this study aligns with the latter. See other recent and different applications to braneworld models in references Nakas et al. (2020, 2024); Neves (2021).
Soon later Hawking et al. demonstrated that the brane geometry can correspond to a four-dimensional Schwarzschild black hole embedded in a five-dimensional AdS spacetime, also known as the black strings (BS) solution Chamblin et al. (2000). It is important to mention that analysis of observations of the shadows of M87* has provided indications that the induced geometry on the brane could represent both ultra-compact objects and wormholes. In particular, reference Banerjee et al. (2022), studying the embedding of such objects, demonstrated that the braneworld scenario fits better with the observed shadow and the diameter of the image of Sgr A* compared to the corresponding situation in general relativity. Even though wormholes, in general, require exotic matter to remain stable, in the presence of extra dimensions such is not the case. Therefore, they showed that the wormhole solution considered does not require exotic matter on the brane. Thus, this reason renewed the interest in embedding other types of geometries different from the Schwarzschild vacuum solution, such as wormholes or regular black holes.
A notable physical consequence is that black string solutions in asymptotically flat space are unstable to wavelength perturbations Gregory and Laflamme (1993). However, as mentioned in reference Chamblin et al. (2000), an AdS spacetime acts as a confining box that prevents the development of fluctuations with wavelengths much greater than $l$, thereby averting this type of instability. Other characteristics of the model discussed in reference Chamblin et al. (2000) include that the Kretschmann invariant diverges at the origin of the radial coordinate and the infinity of the extra dimension (the so-called AdS horizon). Thus, at these two locations, the laws of physics cease to operate.
@@
II Revisinting Hawking et al. model
One way of constructing a black hole solution in a braneworld scenario was proposed by Hawking et al.. in reference Chamblin et al. (2000). The focus of the paper is the AdS singularity. However, since we are also interested in the Black string stability, we will revisit the authors’ results to include this. They began with the Randall-Sundrum metric expressed in terms of the new coordinate $z=le^{y/l}$
$ds^{2}=\frac{l^{2}}{z^{2}}(\eta_{\mu\nu}dx^{\mu}dx^{\nu}+dz^{2}).$ (1)
The authors constructed a black string solution in five dimensions by replacing the Minkowski element in the metric (1) with
$ds^{2}=\frac{l^{2}}{z^{2}}[-U(r)dt^{2}+U(r)^{-1}dr^{2}+r^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2})+dz^{2}],$ (2)
with $U(r)=1-\frac{2M}{r}$, which describes a $4d$ Schwarzschild line element. $M$ is a constant associated with the mass of the four-dimensional black hole embedded into an extra dimension. The coordinates take on values in the following ranges: $t\in(-\infty,+\infty)$, $r\in[0,+\infty)$, $\phi\in[0,2\pi]$, $\theta\in[0,\pi]$, and $z\in(-\infty,+\infty)$. Soon latter, it was shown that the Hawking is unstable under linear perturbations Gregory (2000).
An interesting aspect of this geometry is that it exhibits two singularities: one located along $r=0$, akin to a Schwarzschild black hole, and another singularity at $z=\infty$. These characteristics can be demonstrated by analyzing the Kretschmann scalar, which is given by:
$R_{MNPQ}R^{MNPQ}=\frac{1}{l^{4}}\left(40+\frac{48M^{2}z^{4}}{r^{6}}\right).$ (3)
It’s easy to see that $R_{MNPQ}R^{MNPQ}$ diverges at the AdS horizon ($z=\infty$) and the singularity of the black hole ($r=0$). If one persists in demanding the Schwarzschild metric on the brane, albeit with a standard AdS horizon, the consequence is the necessity of accommodating matter within the bulk Kanti and Tamvakis (2002).
Next, to analyze this singularity, the authors consider the behavior of geodesics. By considering the where affine parameter $\lambda$, the geodesic equation for the $z$ coordinate ($\theta=\pi/2$) is given by
where $\sigma=0,1$ are for null and time-like geodesics. We are interested in solutions in which $z$ can reach the singularity. For null and time-like geodesics, they are given by
$z=-\frac{z_{1}l}{\lambda},z=-z_{1}\csc(\lambda/l).$ (5)
In both cases, $z_{1}$ is a constant fixed so that $\lim_{\lambda\to 0^{-}}z\to\infty$. This can be used to analyze the radial direction. Next, the authors introduce the parameters
$\nu=-\frac{z_{1}^{2}}{\lambda},\;\nu=-(z_{1}^{2}/l)\cot(\lambda/l)$ (6)
$r=z_{1}r/l,\;\bar{E}=z_{1}E/l,\;\bar{L}=z_{1}^{2}L/l^{2},\;\bar{M}=z_{1}M/l.$ (7)
This leaves the radial equation in the form
$\left(\frac{d\bar{r}}{d\nu}\right)^{2}+\left(1+\frac{\bar{L}^{2}}{\bar{r}^{2}}\right)\left(1-\frac{2\bar{M}}{\bar{r}}\right)=\bar{E}^{2}.$ (8)
where $\bar{M}$ is the ADM mass for an observer at $z=z_{0}=l^{2}/z_{1}$. With the above equation, the authors conclude that bound geodesics are divergent at the AdS boundary, but the unbounded ones are not.
In order to conclude about singularity, they calculate the Riemann scalar in a reference frame that propagates parallel to the geodesics. Let the tangent vector to these geodesics with $L=0$ be given by
$u^{M}=\left(\frac{Ez^{2}}{U(r)l^{2}},\frac{z^{2}}{l^{2}}\sqrt{E^{2}-\frac{l^{2}}{z_{1}^{2}}U(r)},0,0,\frac{z}{l}\sqrt{\frac{z^{2}}{z_{1}^{2}}-1}\right),$ (9)
The unit normal vector to the geodesic $n^{M}$, which is parallel propagated along it, is given by
$n^{N}=\left(-\frac{zz_{1}}{l^{2}U(r)}\sqrt{E^{2}-\frac{l^{2}}{z_{1}^{2}}U(r)},-\frac{Ez_{1}z}{l^{2}},0,0,0\right).$ (10)
When is straightforward to check that $u^{P}\nabla_{P}n^{Q}=0$. Thus, by calculating one of the curvature components in this reference frame, we obtain
$R_{(u)(n)(u)(n)}=R_{MNPQ}u^{M}n^{N}u^{P}n^{Q}=\frac{1}{l^{2}}\left(1-\frac{z^{4}}{z_{1}^{2}}\frac{2M}{r^{3}}\right),$ (11)
which diverges along the geodesic to $\lambda\to 0^{-}$. The same conclusion is obtained considing $\bar{L}=0$ in Eq.(8). We get
$\left(\frac{d\bar{r}}{d\nu}\right)^{2}+\left(1-\frac{2\bar{M}}{\bar{r}}\right)=\bar{E}^{2},$ (12)
and, as expected, the solution diverges at $\bar{r}=0$.
On the other hand, following the idea of Randall Sundrum, a mass parameter $M$ located on a brane located at $y=y_{0}$ in an effective $4D$ theory is given by $M_{e}=\exp(-y_{0}/l)M$. Following reference Chamblin et al. (2000), under a change of coordinates $w=z-z_{0}$, the metric maintains its structure, and the wall also could be located at $w=0\to z=z_{0}$. Thus, using $z=le^{y/l}$, the mass measured by an inhabitant of the wall at $z=z_{0}$ would be $M_{e}=Ml/z_{0}$. This implies that the proper radius of the horizon in five dimensions is $r_{e}=2M_{e}$. Assuming that the AdS radius $l$ is on the order of the Planck length, it should hold that $M/z_{0}>>1$ to be consistent with astrophysical data.
It is well known that black string solutions in asymptotically flat space are unstable to wavelength perturbations Gregory and Laflamme (1993). In the case of RS black holes, as indicated by reference Chamblin et al. (2000), an AdS spacetime acts like a confining box that prevents fluctuations with wavelengths much greater than $l$ from developing, thus preventing the occurrence of this type of instability. Thus, if the curvature scalar $R=-20/l^{2}$ is very small, the AdS curvature is negligible, and the space-time behaves like a flat space. This means that the black string solution will be unstable to perturbations with wavelengths on the order of the horizon radius $r_{e}=\frac{2Ml}{z_{0}}$ such that $\frac{r_{e}}{l}=\frac{2M_{e}}{l}=\frac{2M}{z_{0}}\ll 1$. In other words, for very large values of $z$, such perturbations will fit into the AdS box, so instability could occur near the AdS horizon. However, for $\frac{2M}{z_{0}}\gg 1$, the potential instability would occur at wavelengths much larger than $l$, i.e., $\frac{r_{e}}{l}=\frac{2M}{z_{0}}\gg 1$, and, therefore, the AdS curvature prevents instability. Otherwise, an instability can occur, but if it does, it must do so at smaller wavelengths. Thus, the black string solution is stable far from the AdS horizon but potentially unstable near it. The question was settled when it was shown that RS black holes are unstable under linear perturbations Gregory (2000).
A simple way to study instability is by the use of squared sound speed Capozziello and Godani (2022) along a direction (coordinate) $i$. It is given by
$(v^{i}_{s})^{2}=\frac{d<p>}{d\rho},$ (13)
where $<p>$ is the average pressure. To ensure stability in the solution, the speed of sound must satisfy $(v^{i}_{s})^{2}\geq 0$, with an additional constraint of $v^{i}_{s}<1$ in order to be physical. We, therefore, conclude that the solution is unstable for the Hawing model since $d<p>/d\rho=-1$. We will use this method to study the instability of the black-bounce RS black hole.
III Simpson-Visser regularization method applied to Hawking model
To eliminate the radial singularity of equations (3), (11) and (12), we can consider the Simpson-Visser regularization Simpson and Visser (2019) on the Schwarzschild line element 2. Therefore, now consider the metric given by
@@ III Simpson-Visser regularization method applied to Hawking model
We can analyze the causal structure of this geometry by taking $\theta=\pi/2$, $\phi=0$, $z=cte=z_{0}$, and $ds=0$, from which we obtain the following expression for $dr/dt$
@@
III.1 Ricci and Kretschmann invariants
@@
It is important to note that, as expected, taking $a\to 0$ returns (3). Here we assume the general case where $a\neq 0$, so that we have the following asymptotic behaviors
@@
From equations (22), (23), we can deduce that the spacetime $5D$ is regular at the origin $r=0$ (or $\tilde{r}=a$), unlike (3), which is singular at the origin. On the other hand, from equations (24), (25), we observe that at the infinite of the radial coordinate, the values of the scalars Ricci and Kretschmann corresponding to an AdS space are recovered. Thus, unlike Chamblin et al. (2000), which has a global AdS behavior, in our case such behavior only occurs locally at the radial coordinate’s asymptote.
In addition, $\mathcal{K}$ continues to diverge in $z\to\infty$, which suggests that the singularity at the AdS horizon remains. It is important to mention that in the case where the extra coordinate is compactified as $\bar{z}=\bar{R}z$, where $\bar{R}$ is the compactification radius, the new domain would be $\bar{z}\in[-\pi,\pi]$, thus suppressing the singularity at the AdS horizon. In this context, the $\mathcal{Z}_{2}$ symmetry induces the presence of Dirac delta functions, located at the brane locations, in the equations of motion and consequently in the expression of the curvature scalar. This suggests that one interpretation could be that the presence of this Dirac delta distribution in the invariants does not imply a physical singularity like that produced by any diverging function, as for example in eq (3), but represents a pathology associated with mirror symmetry and the confinement of matter. It is worth mentioning that singularities of functions, in general, are not volume integrable, unlike delta distributions, which are.
III.2 Geodesic analysis
In this section, we will carefully analyze the geodesic movement and compare it to the results of section II. We consider the geodesic equation with $\theta=\pi/2$ and study the $z$ and radial directions. For the $z$ coordinate we get
@@ III.2 Geodesic analysis
where $\sigma=0$ for null geodesics and $\sigma=1$ for time-like geodesics. Equation (28) is the same equation obtained by (4), which shows that the regularization applied does not interfere with the movement in the extra direction. It is interesting to point out that this is consistent with the analysis of stability analysis in the next section. Thus, the solutions to the above equation are (5). With the same behavior as before.
@@
By using the definitions (6) and (7) and defining $\bar{a}=z_{1}a/l$, we get the radial equation
@@
This equation is equivalent to the radial equation of time-like geodesics with mass $\bar{M}=Mz_{1}/l$ and regularization parameter $\bar{a}=az_{1}/l$ in a Simpson-Visser like geometry Simpson and Visser (2019). According to Chamblin et al. (2000), $\bar{M}$ is the ADM mass for an observer at $z=z_{0}=l^{2}/z_{1}$. Once again, we can view it as the usual radial geodesic equation $\dot{\bar{r}}^{2}+V_{eff}=\bar{E}^{2}$. For $\bar{L}=0$, the singularity is excluded, in comparison with (12).
@@
which diverges along the geodesic to $\lambda\to 0^{-}$. Hence, our case remains regular at the origin, unlike (11); however, there is still a divergence at the AdS horizon.
IV Energy Conditions and Stability
@@
For $a=0$, we have $\rho_{E}=-p_{1}=-p_{2}=-p_{3}=-6/(l^{2}\kappa_{5}^{2})\equiv\Lambda_{5D}$, which is the cosmological constant term.
It is direct to check that the induced metric on the brane is given by the four-dimensional SV spacetime. The energy-momentum tensor located at the brane location $y=0$ in the original coordinates is given by $T^{\mu}_{\nu}=\delta(y)\text{diag}(\sigma,\sigma,\sigma,\sigma)$, where $\sigma$ is the brane tension. Given the absence of sources or matter fields on the brane, we can assert that the induced SV geometry on the brane should arise from the influence of the geometrical and matter deformations in the bulk. Using the junction conditions, for the line element (16), we can check that the brane tension is given by $\sigma=\dfrac{6}{\kappa^{2}l}>0$.
IV.1 Energy conditions

### v2

@@ Abstract
We provide a way for embedding a 4-dimensional geometry corresponding to the Simpson-Visser (SV) spacetime—which is capable of representing a traversable wormhole, a one-way wormhole, or a regular black hole—into a Randall-Sundrum setup. To achieve this, we linearly deform the bulk geometry and the bulk matter distribution concerning a coupling constant. These deformations induce a transition from a $5D$ vacuum AdS state to an anisotropic matter distribution. The latter results in the induced geometry on the brane transitioning from a singular Schwarzschild spacetime to a regularized SV spacetime. Since there are no sources or matter fields on the brane, we can assert that the induced SV geometry on the brane arises from the influence of geometrical and matter deformations in the bulk. Thus, the central singularity is suppressed. We determine the cases where the energy conditions are either satisfied or violated. Our spacetime is asymptotically radial AdS, which is intriguing given the absence of a global AdS box that would prevent instability under larger wavelength perturbations. Therefore, it is no longer appropriate to claim that instability exists for very small perturbations near the AdS horizon. Thus, we propose that the stability of the solution can be analyzed by examining the speed of sound due to the presence of matter fields in the energy momentum tensor.
@@ I Introduction
The RS model involves the embedding of $4D$ flat branes, i.e., whose geometry corresponds to Minkowski spacetime. Since its inception, various generalizations of the Randall-Sundrum model, featuring different types of four-dimensional geometries including black holes and wormholes, have emerged. The approach to studying these generalizations is primarily based on two main approximations: one originates from the brane, and the other from the bulk. In the former, the four-dimensional black holes or wormholes are constructed by studying the equations of motion induced on the brane Shiromizu et al. (2000). Some examples in references Molina and Neves (2012); Neves (2015); Neves and Molina (2012); Casadio et al. (2015). However, this methodology does not fully allow for testing the influence of bulk geometry on the physical and geometric behavior of the four-dimensional geometry. Regarding the latter, the second approach constructs spacetimes directly from the bulk, and the brane equations are derived from there. See Chamblin et al. (2000); Kanti et al. (2002); Creek et al. (2006). The approach embraced in this study aligns with the latter. See other recent and different applications to braneworld models in references Nakas et al. (2020, 2024); Neves (2021).
It is important to mention that analysis of observations of the shadows of M87* has provided indications that the induced geometry on the brane could represent both ultra-compact objects and wormholes. In particular, reference Banerjee et al. (2022), studying the embedding of such objects, demonstrated that the braneworld scenario fits better with the observed shadow and the diameter of the image of Sgr A* compared to the corresponding situation in general relativity. Even though wormholes, in general, require exotic matter to remain stable, in the presence of extra dimensions such is not the case. Therefore, they showed that the wormhole solution considered does not require exotic matter on the brane. Thus, this reason renewed the interest in embedding other types of geometries different from the Schwarzschild vacuum solution, such as wormholes or regular black holes.
@@
II Simpson-Visser regularization
In order to represent the embedding of a singularity-free object in a braneworld setup, we can consider that this object corresponds to what results from applying the Simpson-Visser regularization Simpson and Visser (2019) to a $4D$ Schwarzschild space. Thus, we will write the $5D$ line element as:
@@ II Simpson-Visser regularization
We can analyze the causal structure of this geometry by taking $\theta=\pi/2$, $\phi=0$, $z=cte=z_{0}$, and $ds^{2}=0$, from which we obtain the following expression for $dr/dt$
@@
II.1 Ricci and Kretschmann invariants
@@
It is important to note that, as expected, taking $a\to 0$ causes the central singularity at $r\to 0$ to reappear. Here we assume the general case where $a\neq 0$, so that we have the following asymptotic behaviors
@@
From equations (9), (10), we can deduce that the spacetime $5D$ is regular at the origin $r=0$ (or $\tilde{r}=a$). On the other hand, from equations (11) and (12), we observe that at infinity in the radial coordinate, the Ricci and Kretschmann scalars recover the values corresponding to an AdS space. Thus, our solution only behaves locally as AdS at the boundary of the asymptotic radial region.
In addition, we observe that $\mathcal{K}$ diverges as $z\to\infty$, which suggests the presence of a singularity at what is referred to as the AdS horizon. It is important to mention that in the case where the extra coordinate is compactified as $\bar{z}=\bar{R}z$, where $\bar{R}$ is the compactification radius, the new domain would be $\bar{z}\in[-\pi,\pi]$, thus suppressing the singularity at the AdS horizon. In this context, the $\mathcal{Z}_{2}$ symmetry induces the presence of Dirac delta functions, located at the brane locations, in the equations of motion and consequently in the expression of the curvature scalar. This suggests that one interpretation could be that the presence of this Dirac delta distribution in the invariants does not imply a physical singularity like that produced by any diverging function, but represents a pathology associated with mirror symmetry and the confinement of matter. It is worth mentioning that singularities of functions, in general, are not volume integrable, unlike delta distributions, which are.
II.2 Geodesic analysis
In this section, we will carefully analyze the geodesic movement. We consider the geodesic equation with $\theta=\pi/2$ and study the $z$ and radial directions. For the $z$ coordinate we get
@@ II.2 Geodesic analysis
where $\sigma=0$ for null geodesics and $\sigma=1$ for time-like geodesics.
For null and time-like geodesics, the solutions of this latter equation are given by
$z=-\frac{z_{1}l}{\lambda},z=-z_{1}\csc(\lambda/l).$ (16)
In both cases, $z_{1}$ is a fixed constant such that $\lim_{\lambda\to 0^{-}}z\to\infty$. Thus, we can deduce that even with the radial regularization of Simpson-Visser, an object would be destroyed upon reaching the infinity of the radial coordinate. It is interesting to point out that this is consistent with the analysis of stability analysis in the next section.
@@
By using the definitions
$\nu=-\frac{z_{1}^{2}}{\lambda},\;\nu=-(z_{1}^{2}/l)\cot(\lambda/l)$ (20)
$r=z_{1}r/l,\;\bar{E}=z_{1}E/l,\;\bar{L}=z_{1}^{2}L/l^{2},\;\bar{M}=z_{1}M/l.$ (21)
and defining $\bar{a}=z_{1}a/l$, we get the radial equation
@@
This equation is equivalent to the radial equation of time-like geodesics with mass $\bar{M}=Mz_{1}/l$ and regularization parameter $\bar{a}=az_{1}/l$ in a Simpson-Visser like geometry Simpson and Visser (2019). According to Chamblin et al. (2000), $\bar{M}$ is the ADM mass for an observer at $z=z_{0}=l^{2}/z_{1}$. Once again, we can view it as the usual radial geodesic equation $\dot{\bar{r}}^{2}+V_{eff}=\bar{E}^{2}$. For $\bar{L}=0$, the singularity is excluded, in comparison with the case where the embedded $4D$ geometry corresponds to that of Schwarzschild.
@@
which diverges along the geodesic to $\lambda\to 0^{-}$. Hence, our case remains regular at the origin; however, there is still a divergence at the AdS horizon. However, the presence of a singularity at infinity in the radial coordinate could make sense from a physical perspective by considering the behavior of the geometry at the radial boundary. At this boundary, our solution behaves like an AdS space. On the other hand, it is well-known that the AdS radius acts as a sort of box, leading to instabilities near the infinity of the extra coordinate Gregory and Laflamme (1993); Chamblin et al. (2000).
III Energy Conditions and Stability
@@
For $a=0$, we have $\rho_{E}=-p_{1}=-p_{2}=-p_{3}=-6/(l^{2}\kappa_{5}^{2})\equiv\Lambda_{5D}$, which is the negative cosmological constant term.
It is direct to check that the induced metric on the brane is given by the four-dimensional SV spacetime. The energy-momentum tensor located at the brane location $y=0$ in the original coordinates is given by $T^{\mu}_{\nu}=\delta(y)\text{diag}(\sigma,\sigma,\sigma,\sigma)$, where $\sigma$ is the brane tension. Given the absence of sources or matter fields on the brane, we can assert that the induced Simpson-Visser (SV) geometry on the brane should result from the influence of geometrical and matter deformations in the bulk. Therefore, it is not necessary to include additional matter fields at the location of the brane beyond its own tension. Using the junction conditions, for the line element (3), we can check that the brane tension is given by $\sigma=\dfrac{6}{\kappa^{2}l}>0$.
III.1 Energy conditions
First of all, we note that we will use the following definitions to analyze the energy conditions:
Null Energy Condition (NEC): $\,\rho+p_{1}\geq 0,\rho+p_{2}\geq 0,\rho+p_{3}\geq 0$ (27)
Weak Energy Condition (WEC): $\,\rho\geq 0,\rho+p_{1}\geq 0,\rho+p_{2}\geq 0,\rho+p_{3}\geq 0$ (28)
Strong Energy Condition (SEC): $\rho+p_{1}\geq 0,\rho+p_{2}\geq 0,\rho+p_{3}\geq 0,$
$\rho+p_{1}+2p_{2}+p_{3}\geq 0$ (29)
With the expressions of the energy density and pressures obtained, we now examine if the energy conditions are globally satisfied in the bulk.
III.1.1 Energy Density
In Fig.1, we present the plot of the energy density as a function of $r$ for a fixed point in the bulk. Analyzing the graph, it is evident that $\rho_{E}$ is globally negative for a fixed and finite $z$, thereby violating the Weak Energy Condition (WEC) in all three cases under consideration for this region. However, for a fixed $r$, the energy density can take positive values. In fact, for $r$ values close to $r=0$, the energy density is positive for some values of $z$ such that $|z|>z_{1}$.
III.1.2 NEC
$\rho_{E}+p_{1}$ $=$ $-\frac{2a^{2}z^{2}}{l^{2}\kappa_{5}^{2}}\left(\frac{(r^{2}+a^{2})^{1/2}-2M}{(r^{2}+a^{2})^{5/2}}\right),$ (30)
$\rho_{E}+p_{2}$ $=$ $\frac{3a^{2}z^{2}}{l^{2}\kappa_{5}^{2}}\cdot\frac{M}{(r^{2}+a^{2})^{5/2}},$ (31)
From the above expressions, it is clear that the conditions $\rho_{E}+p_{2}\geq 0$ and $\rho_{E}+p_{3}\geq 0$ are satisfied for all values of $r$ and $z$. In Fig. 2, we plot $\rho_{E}+p_{1}$ as a function of $r$ for a fixed value of $z$. It is evident from this plot that the NEC is satisfied only within a specific region for the RBH case and at the asymptotic radial limit of $\rho_{E}+p_{1}$. Therefore, according to Equation (30), the NEC is satisfied in this case only for $r\leq\sqrt{(2M)^{2}-a^{2}}$, which corresponds to $r\leq r_{h}$ for the RBH case and in the limit as $r\to\infty$.
III.1.3 SEC
From the statement of this subsection, we can observe that for the Strong Energy Condition (SEC) to be satisfied, both the Null Energy Condition (NEC) and the following additional condition must be met:
$\rho_{E}+\sum_{i}p_{i}=\frac{1}{l^{2}\kappa_{5}^{2}}\left(18+\frac{a^{2}z^{2}}{\tilde{r}^{5}}(\tilde{r}-M)\right)\geq 0$ (33)
In the previous section, it was determined that the Null Energy Condition (NEC) is satisfied only for $\tilde{r}\leq 2M$ in the RBH case and in the limit as $r\to\infty$. On the Figure 3, we observe that the condition (33) can be satisfied for $a<2M$, corresponding to the RBH case. More rigorously, from Equations (30) and (33), the Strong Energy Condition (SEC) at $a<2M$ is satisfied at for the cases where $\tilde{r}\to\infty$ and where: